# Standard modules of affine Hecke algebras

### v1

@@ Abstract
An unpublished theorem of Clozel, proven with global techniques, says that the class of essentially square-integrable representations is also defined over $\bar{\mathbb{Q}}_{\ell}$. As an application of our main result, we give a local proof of this theorem for $G=\mathrm{GL}_{n}$.
Keywords— Standard representation, essentially square-integrable representation, Iwahori-Hecke algebra, Chow group
@@ 1.1 Representations over $\bar{\mathbb{Q}}_{\ell}$
Let $G$ be a connected reductive group defined and split over a non-archimedean local field $F$. Let $\pi$ be a complex admissible representation of $G(F)$, let $\gamma$ be a field automorphism of $\mathbb{C}$, and define a representation $\gamma\cdot\pi$ of $G(F)$ by declaring that its matrix coefficients are of the form $\gamma(f(g))$, for matrix coefficients $f$ of $\pi$ (see Section 2.2 for a precise definition). It is clear that $\gamma\cdot\pi$ is irreducible when $\pi$ is, and likewise for supercuspidality. In [KSV] and [SolleveldStandard], less obvious heritability from $\pi$ to $\gamma\cdot\pi$ is studied. Recall that the standard representations of $G(F)$ are by definition parabolic inductions $i_{P}^{G}(\sigma\otimes\nu)$ for irreducible tempered representations $\sigma$ and positive unramified characters $\nu$ of of $M_{P}(F)$, respectively, where $M_{P}$ is the Levi subgroup of $P$ (see [SolleveldStandard, Def. 4.1, Thm. 4.3 (a)]). Thus the standard modules occur in algebraic families.
@@
The first result of this note is a new geometric proof, for representations in the principal block of split groups, of Theorem 1. Precisely, we give a geometric proof, as our Corollary 1, that if $K(s,N,\rho)$ is a standard module for the Iwahori-Hecke algebra $H$ of $G$, as constructed in terms of $G^{\vee}$ and the enhanced $L$-parameter $(s,N,\rho)$ in [KLDeligneLanglands], [CG], then $\gamma\cdot K(s,N,\rho)\simeq K(\gamma(s),N,\rho_{\gamma})$, as one would expect upon considering central characters. Here, $\rho_{\gamma}$ is the pullback of $\rho$ along the isomorphism $\pi_{0}(Z_{G^{\vee}}(s,N))\simeq\pi_{0}(Z_{G^{\vee}}(\gamma(s),N))$ induced by $\gamma$, see (10) below.
Note that if $\pi$ is generated by its Iwahori-fixed vectors for a chosen Iwahori subgroup $I$ of $G(F)$, then so is $\gamma\cdot\pi$. In general, the functor $\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}$- does not preserve Bernstein blocks (consider $G=\mathrm{GL}_{1}$), but as it is an auto-equivalence of the category of smooth representations, it still interacts well enough with types and Hecke algebras for our purposes, see Section 3.1.
The main input to our argument is the main result of [DLP], which with [KLDeligneLanglands], [CG] implies that, up the central character, representations $K(s,N,\rho)$ can be constructed on rationalized Chow groups of certain varieties defined over $\mathbb{Q}$. This may be viewed as further suggestion that (at least for the principal block) the standard modules admit a natural interpretation in terms of [FS].
@@
$W_{F}\times\mathrm{SL}_{2}\to G^{\vee}(\mathbb{C})$
@@
Combining Corollary 1 with the results of [BushKutzBook], [BushKutzComp], [BHK], we obtain the second result of this note: We give a local proof, valid for function fields, for $G=\mathrm{GL}_{n}$, of
@@
Finally, note that irreducible tempered and discrete series representations are standard, but that within the class of standard representations, the tempered representations are not stable under $\mathrm{Aut}(\mathbb{C})$. Thus the standard representations are a class of representations defined over $\bar{\mathbb{Q}}_{\ell}$ containing the tempered representations. Our proof of Theorem 1 will show that the standard representations form essentially the minimal enlargement of the tempered representations, at least in the case of the principal block.
In [BravermanKazhdan], Braverman-Kazhdan introduce a certain subalgebra $\mathcal{J}$ of the Harish-Chandra Schwartz algebra of $G(F)$. By [suzuki], Theorem 1 implies that the class of simple $\mathcal{J}$-modules is defined over $\bar{\mathbb{Q}}_{\ell}$. Our proof in this note provides an independent argument for the principal block.
@@ 1.2 Acknowledgements
The author thanks Anne-Marie Aubert, Dinakar Muthiah, Evgeny Shinder, Maarten Solleveld, Eric Sommers, and Yakov Varshavsky for helpful conversations. Some of this work was completed during “Geometric representation theory in the Langlands program” at CIRM Luminy; the author thanks the organizers and staff for their hospitality. The author was supported by Engineering and Physical Sciences Research Council grant UKRI167 “Geometry of double loop groups.”
@@ 2.1 Notation and conventions
Let $G$ be the group dual to $G^{\vee}$, considered as a group over $\mathbb{C}$, with root datum $(X^{*},\Phi,X_{*},\Phi^{\vee})$, and chosen set $\Sigma$ of simple roots. We write $\mathcal{B}$ for the flag variety of $G$, $\tilde{\mathcal{N}}=T^{*}\mathcal{B}$, and $\mathrm{St}=\tilde{\mathcal{N}}\times_{\mathcal{N}}\tilde{\mathcal{N}}$ for the Steinberg variety of $G$, where $\mathcal{N}\subset\mathfrak{g}=\operatorname{\mathrm{Lie}}(G)$ is the nilpotent cone. All of these depend only on $G_{\mathrm{ad}}=G/Z(G)$, a semisimple group of adjoint type. Therefore they are each defined over $\mathbb{Q}$, by Chevalley’s theorem (in fact, over $\mathbb{Z}$). Given a nilpotent element $N\in\mathfrak{g}$, we write $\mathcal{B}_{N}$ for the Springer fibre, and given $g\in G$, we write $X^{g}$ for the fixed points for any $G$-scheme $X$. We will view the above $G$-schemes as $G\times{\mathbb{G}_{\mathrm{m}}}$-scheme, such that $\mathbb{G}_{\mathrm{m}}$} scales the cotangent fibres in $\tilde{\mathcal{N}}$.
@@
[…]
@@ 2.2 Action of field automorphisms
Now we fix once and for all a field automorphism $\gamma\in\mathrm{Aut}(\mathbb{C})$. Following [KSV, §2], [SolleveldStandard, §5], let $\mathbb{C}_{\gamma}$ denote the $\mathbb{C},\mathbb{C}$-bimodule $\mathbb{C}$ with actions
@@
Let $\pi$ be an admissible $G^{\vee}(F)$-module, and let $\gamma\cdot\pi=\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\pi$ be the twist of $\pi$ by $\gamma$. It has scalar multiplication given by
@@
$\left\langle 1\otimes\tilde{v},(\gamma\cdot\pi)(g)(1\otimes v)\right\rangle_{\gamma\cdot\pi}=\left\langle 1\otimes\tilde{v},1\otimes\pi(g)v\right\rangle=\gamma(c_{\alpha})=\gamma\left(\left\langle\tilde{v},\pi(g)v\right\rangle_{\pi}\right).$
Given $f\in\mathcal{H}=C_{c}^{\infty}(G^{\vee}(F))$, then matrix coefficients of the operator $(\gamma\cdot\pi)(f)$ are then by definition the numbers
@@
As we normalize the Haar measure such that open compact subgroups of $G$ have rational volumes, if $f\in C_{c}^{\infty}(G,\mathbb{Q})$ is rational-valued, then the above is equal to
$\gamma\left(\left\langle\tilde{v},\pi(f)v\right\rangle\right),~~~f\in C_{c}^{\infty}(G^{\vee}(F),\mathbb{Q}).$
@@ 2.3 Iwahori-spherical standard modules as vector spaces
For any $(s,N)$ such that $\mathrm{Ad}(s)N=qN$, we have an isomorphism of Chow group and homology
@@ 2.4 Standard modules as modules over the Bernstein subalgebra
We define an action of $G$-equivariant line bundles on the diagonal $\tilde{\mathcal{N}}\subset\mathrm{St}$ on $A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}$: Let $\mathcal{L}$ be such a line bundle, and consider its class as belonging to $K(\mathcal{B})$ via the Thom isomorphism. Consider the natural embeddings
@@
$=\operatorname{Spec}(\gamma)^{*}(\mathrm{ch}_{*}(\iota^{*}\mathcal{L}\otimes\mathcal{E})$ (5)
@@
where between (4) and (5) we used smoothness of $\mathcal{B}^{s}$ and [Fulton, Prop. 18.1 (c)], and likewise between (7) and (8); between (5) and (6) and (7) and (8), we also used that the homological Chern character commutes with pullback by smooth proper morphisms [BFMIHES, (2.6)], [Fulton, Thm. 18.1]. Between (7) and 8 we used that $\operatorname{Spec}(\gamma)^{*}[\mathcal{L}]=[\mathcal{L}]$ in the Grothendieck group, as the associated bundle construction defining each generator $\mathcal{O}_{\lambda}$ of $K(\tilde{\mathcal{N}}/G\times{\mathbb{G}_{\mathrm{m}}})$, $\lambda\in X^{*}(T)$, is defined over $\mathbb{Q}$ (in particular this implies the equality in equivariant $K$-theory, see also Step 1 of the proof of Proposition 1). Between (5) and (6) we used that $\operatorname{Spec}(\gamma)$ is flat and $\iota^{*}\mathcal{L}$ is a line bundle. ∎
@@
The map $\Phi(\gamma)$ intertwines, for each $\mathcal{O}\lambda$ the actions
$\mathcal{O}\lambda(z_{1}\otimes(z_{2}\otimes v))=z_{1}\otimes(\mathcal{O}_{\lambda}\star(z_{2}\otimes v))$
[…]
@@
as required. The Lemma is true for general homology classes by linearity. ∎

### v2

@@ Abstract
An unpublished theorem of Clozel, proven with global techniques, says that the class of essentially square-integrable representations is also defined over $\bar{\mathbb{Q}}_{\ell}$. As an application of our main result, we give a local proof of this theorem for inner forms of $\mathrm{GL}_{n}$, as well as showing that standard representations of these groups are defined over $\bar{\mathbb{Q}}_{\ell}$.
Keywords— Standard representation, essentially square-integrable representation, simple type, Iwahori-Hecke algebra, Chow group
@@ 1.1 Representations over $\bar{\mathbb{Q}}_{\ell}$
Let $G$ be a connected reductive group defined and split over a non-archimedean local field $F$. Let $\pi$ be a complex admissible representation of $G(F)$, let $\gamma$ be a field automorphism of $\mathbb{C}$, and define a representation $\gamma\cdot\pi$ of $G(F)$ by declaring that its matrix coefficients are of the form $\gamma(f(g))$, for matrix coefficients $f$ of $\pi$ (see Section 2.2 for a precise definition). It is clear that $\gamma\cdot\pi$ is irreducible when $\pi$ is, and likewise for supercuspidality. In [KSV] and [SolleveldStandard], less obvious heritability from $\pi$ to $\gamma\cdot\pi$ is studied. Recall that the standard representations of $G(F)$ are by definition parabolic inductions $i_{P}^{G}(\sigma\otimes\nu)$ for irreducible tempered representations $\sigma$ and positive unramified characters $\nu$ of $M_{P}(F)$, respectively, where $M_{P}$ is the Levi subgroup of $P$ (see [SolleveldStandard, Def. 4.1, Thm. 4.3 (a)]). Thus the standard modules occur in algebraic families.
@@
The first result of this note is a new geometric proof, for representations in the principal block of split groups, of Theorem 1. Precisely, we give a geometric proof, as our Corollary 1, that if $K(s,N,\rho)$ is a standard module for the Iwahori-Hecke algebra $H$ of $G$, as constructed in terms of $G^{\vee}$ and the enhanced $L$-parameter $(s,N,\rho)$ in [KLDeligneLanglands], [CG], then $\gamma\cdot K(s,N,\rho)\simeq K(\gamma(s),N,\rho_{\gamma})$, as one would expect upon considering central characters. Here, $\rho_{\gamma}$ is the pullback of $\rho$ along the isomorphism $\pi_{0}(Z_{G^{\vee}}(s,N))\simeq\pi_{0}(Z_{G^{\vee}}(\gamma(s),N))$ induced by $\gamma$, see (10) below. This implies that the Deligne-Langlands correspondence of [KLDeligneLanglands] is defined over $\bar{\mathbb{Q}}_{\ell}$, see Corollary 2.
Note that if $\pi$ is generated by its Iwahori-fixed vectors for a chosen Iwahori subgroup $I$ of $G(F)$, then so is $\gamma\cdot\pi$. In general, the functor $\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}$- does not preserve Bernstein blocks (consider $G=\mathrm{GL}_{1}$), but still interacts well enough with types and Hecke algebras for our purposes, see Lemma 4.
The main input to our argument is the main result of [DLP], which with [KLDeligneLanglands], [CG] implies that, up the central character, representations $K(s,N,\rho)$ can be constructed on rationalized Chow groups of certain varieties defined over $\mathbb{Q}$. This may be viewed as further suggestion that (at least for the principal block or $G=\mathrm{GL}_{n}$) the standard modules admit a natural interpretation in terms of [FS].
@@
$W_{F}\times\mathrm{SL}_{2}(\mathbb{C})\to G^{\vee}(\mathbb{C})$
@@
Combining Corollary 1 with the results of [BushKutzBook], [BushKutzComp], [Secherre], [SecherreStevens], [SecherreStevensIV], we obtain the second result of this note: We give a local proof, valid for function fields, for inner forms $G$ of $\mathrm{GL}_{n}$, of
@@
In fact, using [SolleveldCompletion], we treat general standard modules, giving a proof of Theorem 1 for inner forms of $\mathrm{GL}_{n}$, compatible with 2.
@@
Finally, note that irreducible tempered and discrete series representations are standard, but that within the class of standard representations, the tempered representations are not stable under $\mathrm{Aut}(\mathbb{C})$. Thus the standard representations are a class of representations defined over $\bar{\mathbb{Q}}_{\ell}$ containing the tempered representations. Our proof of Theorem 1 will show that the standard representations form essentially the minimal enlargement of the tempered representations, at least in the case of the principal block and inner forms of $\mathrm{GL}_{n}$.
In [BravermanKazhdan], Braverman-Kazhdan introduce a certain subalgebra $\mathcal{J}$ of the Harish-Chandra Schwartz algebra of $G(F)$. By [suzuki], Theorem 1 implies that the class of simple $\mathcal{J}$-modules is defined over $\bar{\mathbb{Q}}_{\ell}$. Our proof in this note provides an independent argument for the principal block, and for inner forms of $\mathrm{GL}_{n}$. The idempotented piece $\mathcal{J}_{I}$ of $\mathcal{J}$ corresponding to the principal block is isomorphic to Lusztig’s asymptotic Hecke algebra $J$ [BravermanKazhdan] (see the proofs in [Plancherel], [rigid], [BKK]), and [Plancherel] then implies that $\mathcal{J}_{I}$ has a $\mathbb{C}$-basis of $\mathbb{Q}$-valued functions.
1.2 Prior work for discrete series of inner forms of $\mathrm{GL}_{n}$
Theorem 2 can be deduced from a local Langlands correspondence for $G$ satisfying the properties of [KalethaTaibi, Conj. 6.1], see Proposition 6.3 of op. cit.. Of course, this is an extreme approach: all the known proofs [LRS], [Siyan] (for function fields), [Henniart], [HT], [Scholze] (for $p$-adic fields) of the local Langlands correspondence for $\mathrm{GL}_{n}$ and its inner forms use far more global input than does Clozel’s proof.
However, there are also prior local proofs of Theorem 2 for $\mathrm{GL}_{n}$ and its inner forms. After finishing a first version of this paper, we happened upon [Dat, §2.1.1 (ii)], where it is explained that for $\mathrm{GL}_{n}(F)$, Theorem 2 is known to follow the Bernstein-Zelevinski classification, specifically [ZII, Thm. 9.3], or alternatively from [Rogawski] for $p$-adic fields (which again uses global techniques). Presumably the case of $\mathrm{GL}_{n}(D)$ then follows from the (purely local) proof of a Bernstein-Zelevinski-style classification for these groups given by Mínguez-Sécherre in [MSBanal]. As pointed out in [MSJLmodulol, Rem. 10.1], $\mathrm{Aut}(\mathbb{C})$-equivariance of the Jacquet-Langlands correspondence follows from its character-theoretic characterization.
As [BushKutzBook], [BushKutzComp] and [Secherre], [SecherreStevens] are independent of [ZII] and [MSBanal], respectively, our proof is independent of the above-recalled local proofs. The question of standard representations is not taken up in the above works.
1.3 Other groups
In principle, our approach to Theorem 2 based on affine Hecke algebras from types should adapt more readily to general groups than approaches based on the Bernstein-Zelevinski classification. However, the techniques of the present note work only for equal-parameter affine Hecke algebras. While such algebras can appear for types beyond the principal block or inner forms of $\mathrm{GL}_{n}$, it seems that the only groups that can be dealt with via types using equal-parameter algebras exclusively are again inner forms of $\mathrm{GL}_{n}$. Using more general Morita equivalences, Aubert-Baum-Plymen-Solleveld showed that all Bernstein blocks of inner forms of $\mathrm{SL}_{n}$ may be treated by more complicated algebras built out of equal-parameter affine Hecke algebras [ABPS1], but Theorem 7 connecting preservation of standardness to our geometric statement does not seem to be available for these algebras, see the discussion in [SolleveldCompletion, p.43].
@@ 1.4 Acknowledgements
The author thanks Anne-Marie Aubert, Dinakar Muthiah, Evgeny Shinder, Maarten Solleveld, Eric Sommers, and Yakov Varshavsky for helpful conversations, and Maximilien Mackie for pointing out an imprecision in Remark 1 in an earlier version of this note. Some of this work was completed during “Geometric representation theory in the Langlands program” at CIRM Luminy; the author thanks the organizers and staff for their hospitality. The author was supported by Engineering and Physical Sciences Research Council grant UKRI167 “Geometry of double loop groups.”
@@ 2.1 Notation and conventions
Let $G$ be the group dual to $G^{\vee}$, considered as a group over $\mathbb{C}$, with root datum $(X^{*},\Phi,X_{*},\Phi^{\vee})$, and chosen set $\Sigma$ of simple roots. We write $\mathcal{B}$ for the flag variety of $G$, $\tilde{\mathcal{N}}=T^{*}\mathcal{B}$, and $\mathrm{St}=\tilde{\mathcal{N}}\times_{\mathcal{N}}\tilde{\mathcal{N}}$ for the Steinberg variety of $G$, where $\mathcal{N}\subset\mathfrak{g}=\operatorname{\mathrm{Lie}}(G)$ is the nilpotent cone. All of these depend only on $G_{\mathrm{ad}}=G/Z(G)$, a semisimple group of adjoint type. Therefore they are each defined over $\mathbb{Q}$, by Chevalley’s theorem (in fact, over $\mathbb{Z}$ [LusJAMS2009]). Given a nilpotent element $N\in\mathfrak{g}$, we write $\mathcal{B}_{N}$ for the Springer fibre, and given $g\in G$, we write $X^{g}$ for the fixed points for any $G$-scheme $X$. We will view the above $G$-schemes as $G\times{\mathbb{G}_{\mathrm{m}}}$-schemes, such that $\mathbb{G}_{\mathrm{m}}$} scales the cotangent fibres in $\tilde{\mathcal{N}}$.
@@
Let $W$ be the Weyl group of $G$. Let $H$ be the affine Hecke algebra corresponding to $\tilde{W}=W\ltimes X^{*}$ (recall our conventions on duals) over the formal Laurent polynomial ring $\mathbb{Z}[\mathbf{q}^{1/2},\mathbf{q}^{-1/2}]$. We view $\mathbb{Z}[\mathbf{q}^{1/2},\mathbf{q}^{-1/2}]=K_{0}(\mathrm{pt}/{\mathbb{G}_{\mathrm{m}}})$ and identify
@@ 2.2 Action of field automorphisms
Now we fix once and for all a field automorphism $\gamma\in\mathrm{Aut}(\mathbb{C})$. Following [KSV, §2], [SolleveldStandard, §5], let $\mathbb{C}_{\gamma}$ denote the -$\mathbb{C}\text{-}\mathbb{C}$-bimodule $\mathbb{C}$ with actions
@@
Let $\pi$ be an admissible $G^{\vee}(F)$-representation, and let $\gamma\cdot\pi=\mathbb{C}_{\gamma}\otimes_{\mathbb{C}}\pi$ be the twist of $\pi$ by $\gamma$. It has scalar multiplication given by
@@
$\left\langle 1\otimes\tilde{v},(\gamma\cdot\pi)(g)(1\otimes v)\right\rangle_{\gamma\cdot\pi}=\left\langle 1\otimes\tilde{v},1\otimes\pi(g)v\right\rangle_{\gamma\cdot\pi}=\gamma(c_{\alpha})=\gamma\left(\left\langle\tilde{v},\pi(g)v\right\rangle_{\pi}\right).$
Given $f\in\mathcal{H}=C_{c}^{\infty}(G^{\vee}(F))$, the matrix coefficients of the operator $(\gamma\cdot\pi)(f)$ are then by definition the numbers
@@
As we normalize the Haar measure such that open compact subgroups of $G$ have rational volumes, if $f\in C_{c}^{\infty}(G,\mathbb{Q})$ is rational-valued, then (1) is equal to
$\gamma\left(\left\langle\tilde{v},\pi(f)v\right\rangle_{\pi}\right),~~~f\in C_{c}^{\infty}(G^{\vee}(F),\mathbb{Q}).$
@@ 2.3 Iwahori-spherical standard modules as vector spaces
For any $(s,N)$ such that $\mathrm{Ad}(s)N=qN$, we have an isomorphism of Chow groups and homology
@@ 2.4 Standard modules as modules over the Bernstein subalgebra
We define an action of $G$-equivariant line bundles on the diagonal $\Delta\tilde{\mathcal{N}}\subset\mathrm{St}$ on $A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{Q}}$ based on the digression in [CG, §8.2]: Let $\mathcal{L}$ be such a line bundle, and consider its class as belonging to $K(\mathcal{B})$ via the Thom isomorphism. Consider the natural embeddings
@@
$=\operatorname{Spec}(\gamma)^{*}(\mathrm{ch}_{*}(\iota^{*}\mathcal{L}\otimes\mathcal{E}))$ (5)
@@
where between (4) and (5) we used smoothness of $\mathcal{B}^{s}$ and [Fulton, Prop. 18.1 (c)], and likewise between (7) and (8); between (5) and (6) and (7) and (8), we also used that the homological Chern character commutes with pullback by smooth proper morphisms [BFMIHES, (2.6)], [Fulton, Thm. 18.1]. Between (7) and (8) we used that $\operatorname{Spec}(\gamma)^{*}[\mathcal{L}]=[\mathcal{L}]$ in the Grothendieck group, as the associated bundle construction defining each generator $\mathcal{O}_{\lambda}$ of $K(\tilde{\mathcal{N}}/G\times{\mathbb{G}_{\mathrm{m}}})$, $\lambda\in X^{*}(T)$, is defined over $\mathbb{Q}$ (in particular this implies the equality in equivariant $K$-theory, see also Step 1 of the proof of Proposition 1). Between (5) and (6) we used that $\operatorname{Spec}(\gamma)$ is flat and $\iota^{*}\mathcal{L}$ is a line bundle. ∎
@@
The map $\Phi(\gamma)$ intertwines, for each $\mathcal{O}_{\lambda}$ the actions
$\mathcal{O}_{\lambda}(z_{1}\otimes(z_{2}\otimes v))=z_{1}\otimes(\mathcal{O}_{\lambda}\star(z_{2}\otimes v))$
and the $\star$-action defined above on $A_{*}(\mathcal{B}_{N}^{s})_{\mathbb{C}}$.
@@
as required. The Lemma is true for general cycle classes by linearity. ∎